1 volume of gaseous compound C,H,O on complete combustion in presence of 2.5 volume of O2 gives 2 volume of steam and 2 volume of CO2 ,what is the formula of compound if all measurements are made of STP

Grade:11

1 Answers

Vikas TU
7 years ago
P.O.A.C can be used here as the combustion process is total complete.
Let the hydrocarbon be => CxHyOz.
Eqn. would be:
CxHyOz    +     (2x + y/2 – z)O2  =======>             xCO2                  +           (y/2)H2O
POAC for C we get,
x * 1 = 1*2
x = 2
POAC for H we get,
y*1 = 2*2
y = 4
POAC for O we get,
z*1 + 5*2/2 = 2*2 + 1*2
z = 1
Hence the hydrocarbon is:
C2H4O